OpenStudy (anonymous):

How much compound interest is earned on $11,250 principal, with an annual interest rate of 3% compounded annually, over 4 1/2 years? compound interest: I = P (1 + r)t A. $1518.75 B. $1600.50 C. $12,869.55 D. $12,850.50

OpenStudy (anonymous):

@soprano.h.d0816

OpenStudy (misssmartiez):

Nah Uh, hold it!

OpenStudy (misssmartiez):

That isn't the formula for Compound interests, however, for BALANCE.

OpenStudy (misssmartiez):

Correction: B = P(1+ R)^T Then, after we find the B, or balance, then we do P - B = Compound Interest.

OpenStudy (misssmartiez):

1: We plug it in the format: $11,250(1+0.03 (or 3%) )^4.5 Lets do pemdas!

OpenStudy (misssmartiez):

(1+ 0.03) = 1.03 1.03^4.5 = 1.14226668682 Now, we found out the parenthesis format so, $11,250 x 1.14226668682 = $12850.5002267 Balance: 12850.5002267

OpenStudy (misssmartiez):

We aren't done yet, subtract that by the principal! So, 1600.5002267 (How I got it, 11,250 - $12850.5002267)

OpenStudy (misssmartiez):

Round it, $1,600.50 Your compound interest, my friend, is 1,600.50 dollars. Show your steps!

OpenStudy (misssmartiez):

Undoubtedly, you might have more questions on this, remember the formulas: Simple interest: P x R x T Balance: P(1+r)^t Compound interest: P(1+r)^t = B. B - P = Compound interest. You get it @OpenStudier0 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you so much

OpenStudy (misssmartiez):

:) No problem! Tag me if you need anymore assistance!